how many teams of three can you make from a class of 27?
Answer by Alan3354(69443)

how many teams of three can you make from a class of 27?
-----------
the 1st choice is 1 of 27, then 1/26, then 1/25
= 27*26*25 = 17550
-------------
but choosing A, B & C is the same at B, A & C, so divide by 3*2*1
17550/6 = 2925 ways